# Study of $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$

### v1

@@ V.3 $CP$-violation Asymmetry
[…]
@@ VI Systematic uncertainties
  • Number of $J/\psi$ events: The number of $J/\psi$ events is determined to be $(10087\pm 44)\times 10^{6}$ from counting the inclusive hadronic events; the uncertainty is 0.4% Ablikim et al. (2022).
  • Branching fraction of $J/\psi\rightarrow\gamma\eta$: The uncertainty of 1.2% on the branching fraction of $J/\psi\rightarrow\gamma\eta$ is taken from the PDG Navas et al. (2024).
@@
  • PID: The pion PID efficiency for data agrees within 1.0% with that of the MC simulation in the pion momentum region, as reported in Ablikim et al. (2021). There is no specific decay process available for us to study the PID of muons in the low momentum region. Because the muon mass is similar to the pion mass, 1.0% is taken as the systematic uncertainty for the muon Ablikim et al. (2021). Thus, 4.0% is taken as the systematic uncertainty from PID for the $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ decay.
  • Kinematic fit: To investigate the systematic uncertainty associated with the kinematic fit, the track helix parameter correction method Ablikim et al. (2013) is used. Half of the difference in the detection efficiencies with and without the helix corrections is taken as the systematic uncertainty.
@@
  • Background estimation: The background simulations are based on the amplitude analysis and are relatively precise. Therefore, the systematic uncertainty from background models is neglected. In the fit, the numbers of events for specific backgrounds are fixed according to the branching fractions from the PDG. To estimate the effect of the uncertainties of the used branching fractions, a set of random numbers is generated within the uncertainty range of each branching fraction. Using these random scaling parameters, a series of fits to the invariant mass distributions of $\pi^{+}\pi^{-}e^{+}e^{-}$ are performed. The variance of the determined number of signal events and asymmetry parameters is taken as the systematic uncertainties. In the TFF measurement, we perform alternative fits by changing the branching fraction by $-1.0\sigma$ and $+1.0\sigma$, and the largest difference from the nominal results is taken as the systematic uncertainty.
  • Resolution: To estimate the uncertainty from the resolution, we perform alternative fits by changing the resolution from $-1.0\sigma$ to $+1.0\sigma$ for the $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay. The largest difference from the nominal result is taken as the systematic uncertainty on the asymmetry parameter.
@@
[…]
Number of $J/\psi$ events $0.4$.. ... $0.4$
$\mathcal{B}(J/\psi\rightarrow\gamma\eta)$ $1.2$.. ... $1.2$
MC statistics $0.2$.. ... $0.2$
MDC tracking $4.1$.. $1.0$ $0.7$
Photon detection $0.5$.. ... $0.5$
@@
Combine PID and kinematic fit $3.1$ $0.9$ $0.2$..
Photon conversion rejection $1.1$ $1.6$ $1.6$..
Signal shape $3.2$.. ... ...
Generator model $0.1$.. ... ...
Background estimation $0.5$ $2.2$ $0.0$..
Resolution ... $1.8$.. ...
Width ... ... $0.3$..
@@ VII Upper limit of $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$
$\mathcal{B}<\frac{N^{\mathrm{UP}}}{N_{J/\psi}\cdot\mathcal{B}(J/\psi\rightarrow\gamma\eta)}=4.0\times 10^{-7}.$ (12)
@@ VIII Upper limits for ALPs in $\eta\rightarrow\pi^{+}\pi^{-}a,a\rightarrow e^{+}e^{-}$
To study the unknown ALPs decaying into $e^{+}e^{-}$, we perform 40 simulations under the hypothesis $a\rightarrow e^{+}e^{-}$ by varying masses of $a$ in steps of $5$ MeV/$c^{2}$ from 0 to 200 MeV/$c^{2}$.
@@ IX SUMMARY
With a sample of $(10087\pm 44)\times 10^{6}$ $J/\psi$ events, the decays of $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ ($l=e$ or $\mu$) are studied. For the decay $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$, the branching fraction is determined to be $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})=(3.07\pm 0.12_{\rm{stat.}}\pm 0.19_{\rm{syst.}})\times 10^{-4}$, which is in good agreement with theoretical predictions Petri (2010); Borasoy and Nissler (2007) and previous measurements Berlowski et al. (2008); Akhmetshin et al. (2001); Ambrosino et al. (2009), as shown in Fig. 8.
Figure 8: The branching fraction of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ from different theoretical predictions Petri (2010); Borasoy and Nissler (2007) (blue squares), other experiments Berlowski et al. (2008); Akhmetshin et al. (2001); Ambrosino et al. (2009) (red triangles) and this measurement (green dot).
@@
Finally, ALPs are searched for via the decay $\eta\rightarrow\pi^{+}\pi^{-}a,a\rightarrow e^{+}e^{-}$, and the 90% C.L. upper limits on the branching fraction relative to that of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ are presented as a function of the ALP mass within $5-200\ \mathrm{MeV}/c^{2}$, as shown in Fig. 7. The significance for each case is less than $0.5\sigma$.

### v2

@@ V.3 $CP$-violation Asymmetry
$=\frac{\int^{\pi}_{-\pi}F(\varphi)\,\rm{sign}(\sin{2\varphi})\,d\varphi}{\int^{\pi}_{-\pi}F(\varphi)\,d\varphi}$
@@ VI Systematic uncertainties
  • Number of $J/\psi$ events: The number of $J/\psi$ events is determined to be $(10087\pm 44)\times 10^{6}$ from counting the inclusive hadronic events; the uncertainty is 0.4% Ablikim and others (2022).
  • Branching fraction of $J/\psi\rightarrow\gamma\eta$: The uncertainty of 1.2% on the branching fraction of $J/\psi\rightarrow\gamma\eta$ is taken from the PDG Navas and others (2024).
@@
  • PID: The pion PID efficiency for data agrees within 1.0% with that of the MC simulation in the pion momentum region, as reported in Ablikim and others (2021). There is no specific decay process available for us to study the PID of muons in the low momentum region. Because the muon mass is similar to the pion mass, 1.0% is taken as the systematic uncertainty for the muon Ablikim and others (2021). Thus, 4.0% is taken as the systematic uncertainty from PID for the $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$ decay.
  • Kinematic fit: To investigate the systematic uncertainty associated with the kinematic fit, the track helix parameter correction method Ablikim and others (2013) is used. Half of the difference in the detection efficiencies with and without the helix corrections is taken as the systematic uncertainty.
@@
  • Background estimation: The background simulations are based on the amplitude analysis and are relatively precise. Therefore, the systematic uncertainty from background models is neglected. In the fit, the numbers of events for specific backgrounds are fixed according to the branching fractions from the PDG. To estimate the effect of the uncertainties of the used branching fractions, a set of random numbers is generated within the uncertainty range of each branching fraction. Using these random scaling parameters, a series of fits to the invariant mass distributions of $\pi^{+}\pi^{-}e^{+}e^{-}$ are performed. The variance of the determined number of signal events and asymmetry parameters is taken as the systematic uncertainties. In the TFF measurement, we perform alternative fits by changing the branching fraction by $1.0\sigma$ and $1.0\sigma$, and the largest difference from the nominal results is taken as the systematic uncertainty.
  • Resolution: To estimate the uncertainty from the resolution, we perform alternative fits by changing the resolution from $1.0\sigma$ to $1.0\sigma$ for the $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ decay. The largest difference from the nominal result is taken as the systematic uncertainty on the asymmetry parameter.
@@
$\mathcal{B}$ $\mathcal{A}_{CP}$ $m_{V}$ $\mathcal{B}$
Number of $J/\psi$ events $0.4$ ... ... $0.4$
$\mathcal{B}(J/\psi\rightarrow\gamma\eta)$ $1.2$ ... ... $1.2$
MC statistics $0.2$ ... ... $0.2$
MDC tracking $4.1$ ... $1.0$ $0.7$
Photon detection $0.5$ ... ... $0.5$
@@
Combine PID and kinematic fit $3.1$ $0.9$ $0.2$ ...
Photon conversion rejection $1.1$ $1.6$ $1.6$ ...
Signal shape $3.2$ ... ... ...
Generator model $0.1$ ... ... ...
Background estimation $0.5$ $2.2$ $0.0$ ...
Resolution ... $1.8$ ... ...
Width ... ... $0.3$ ...
@@ VII Upper limit of $\eta\rightarrow\pi^{+}\pi^{-}\mu^{+}\mu^{-}$
[…]
@@ VIII Upper limits for ALPs in $\eta\rightarrow\pi^{+}\pi^{-}a,a\rightarrow e^{+}e^{-}$
[…]
@@ IX SUMMARY
With a sample of $(10087\pm 44)\times 10^{6}$ $J/\psi$ events, the decays of $\eta\rightarrow\pi^{+}\pi^{-}l^{+}l^{-}$ ($l=e$ or $\mu$) are studied. For the decay $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$, the branching fraction is determined to be $\mathcal{B}(\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-})=(3.07\pm 0.12_{\rm{stat.}}\pm 0.19_{\rm{syst.}})\times 10^{-4}$, which is in good agreement with theoretical predictions Petri (2010); Borasoy and Nissler (2007); Zillinger et al. (2022) and previous measurements Berlowski and others (2008); Akhmetshin and others (2001); Ambrosino and others (2009), as shown in Fig. 8.
Figure 8: The branching fraction of $\eta\rightarrow\pi^{+}\pi^{-}e^{+}e^{-}$ from different theoretical predictions Petri (2010); Borasoy and Nissler (2007); Zillinger et al. (2022) (blue squares), other experiments Berlowski and others (2008); Akhmetshin and others (2001); Ambrosino and others (2009) (red triangles) and this measurement (green dot).
@@
[…]